### 10 Mar 2026 v1

Abstract
This paper investigates the dynamics and optimal harvesting of age-structured populations governed by McKendrick–von Foerster equations, contrasting two distinct harvesting mechanisms: rate-control and effort-control. For the rate-control formulation, where harvesting acts as a direct additive removal term, we derive first-order necessary optimality conditions of Pontryagin type for the associated infinite-horizon optimal control problem, explicitly characterizing the adjoint system, transversality conditions, and control switching laws. In contrast, the effort-control formulation introduces harvesting as a multiplicative mortality intensity dependent on aggregate population size. We demonstrate that this aggregate dependence structurally alters the optimality system, formally generating a nonlocal coupling term in the adjoint equation that links all ages through the total stock. By combining rigorous variational control derivations, and explicit stationary profiles, this work clarifies the profound mathematical and bioeconomic distinctions between additive and multiplicative harvesting strategies.
keywords:
Age-structured population dynamics; Optimal harvesting; Infinite-horizon optimal control; McKendrick-von Foerster equation; Pontryagin maximum principle
1 Introduction
In this paper we focus on two harvesting mechanisms in age-structured models. The first is a rate-control formulation, in which harvesting enters the state equation as a direct removal term. The second is an effort-control formulation, in which harvesting acts as an additional mortality intensity and the instantaneous yield is proportional to both effort and stock abundance. First, for the infinite-horizon harvesting problem, we derive necessary optimality conditions of Pontryagin type, including the adjoint system, transversality condition, switching relations for the distributed harvesting control and the boundary inflow control, and complementary slackness for the state constraint. We then compare rate-control and effort-control formulations to highlight the structural differences between them. In particular, we exhibit the nonlocal coupling term in the effort-control adjoint equation induced by aggregate population dependence and thereby make explicit how the choice of harvesting mechanism changes the analytical structure of the optimality system.
Age-structured population models provide a classical framework for describing the evolution of heterogeneous populations in which the demographic state depends explicitly on age. In continuous time, such dynamics are typically formulated through McKendrick–von Foerster type transport equations [mkendrick_applications_1925, von_foerster_remarks_1959], which have become a standard mathematical tool in population dynamics, demography, resource management, and harvesting theory. A systematic mathematical treatment of these models is developed in the foundational works of Webb [webb_semigroup_1984, webb_theory_1985], Iannelli [iannelli_mathematical_1995], Metz et al. [metz_dynamics_1986], and Iannelli and Milner [iannelli_basic_2017]. Within this framework, deriving optimal harvesting policies under dynamical constraints becomes particularly central.
The literature relevant to the present paper falls into several closely related strands: the general theory of structured population dynamics and semigroup formulations [webb_semigroup_1984, webb_theory_1985, iannelli_mathematical_1995, metz_dynamics_1986, iannelli_basic_2017, engel_one-parameter_2000, banasiak_perturbations_2006, cushing_introduction_1998, magal_theory_2018]; steady-state analysis for structured populations [gurtin_non-linear_1974, gurtin_simple_1979, cushing_existence_1984, cushing_introduction_1998, pruss_equilibrium_1981, pruss_stability_1983, busenberg_global_1991, inaba_age-structured_2006, inaba_stable_2017, busenberg_endemic_1988, walker_positive_2009]; optimal control and harvesting of age-structured systems [brokate_pontryagins_1985, barbu_optimal_1999, anita_optimal_1998, anita_analysis_2000, anita_optimal_2009, boucekkine_optimal_2013, zhao_optimal_2005, fister_optimal_2004]; infinite-horizon optimal control and transversality [aseev_pontryagin_2004, aseev_infinite-horizon_2012, aseev_maximum_2015]; and bioeconomic formulations of harvesting [clark_economics_1973, clark_economics_1975, clark_mathematical_2010, tahvonen_economics_2009, getz_population_1989, reed_optimal_1979]. A more detailed review of these contributions is given in Section 2.
The contributions of the paper can be summarized as follows:
  1. (i)
    We derive necessary optimality conditions for an infinite-horizon age-structured harvesting problem. They include the adjoint equation, boundary and transversality conditions, switching conditions for the controls, and complementary slackness relations associated with the nonnegativity constraint on the state.
  2. (ii)
    We clarify the mathematical and economic distinction between rate-control and effort-control mechanisms by comparing the way they enter both the state dynamics and the profit functional, and by exhibiting the nonlocal adjoint term induced by aggregate population dependence in the effort-control setting.
Figure 1 provides a mechanistic map of the paper. The paper is organized as follows. Section 2 reviews the relevant literature on age-structured population dynamics, steady-state analysis, optimal control, and bioeconomic harvesting models. In Section 3 we introduce the models and standing assumptions. In Section 4 we formulate the infinite-horizon optimal control problem and derive necessary optimality conditions. Section 5 discusses the effort-control formulation and the associated formal adjoint equation with nonlocal dependence. Section  6 collects autonomous reductions and explicit stationary formulas, and Section 7 concludes with a discussion of the mathematical implications of the two harvesting mechanisms and possible directions for future work. Appendix A presents numerical simulations.
Common state variables age $a$, time $t$, density $x(t,a)$, inflow $p(t)$ aggregate stock: $\scriptstyle E(t)=\int_{0}^{A}x(t,a)\,da$ Model R: Rate-control $\partial_{t}x+\partial_{a}x=$ $\scriptstyle$- $\mu(t,a)x-u(t,a)$ harvesting enters additively affine transport structure Infinite-horizon optimal harvesting maximize discounted profit $\scriptstyle\int_{0}^{\infty}e^{-rt}\times$ $\scriptstyle\Bigl(\int_{0}^{A}c\,u\,da-kp\Bigr)\,dt$ subject to $x\geq 0$ Model E: Effort-control $\scriptstyle\partial_{t}x+\partial_{a}x=$ $\scriptstyle-(\mu(E(t),a)$ $\scriptstyle+w(t,a))x$ harvesting enters multiplicatively nonlinear + nonlocal dependence Pontryagin system for Model R $\scriptstyle\partial_{t}\lambda+\partial_{a}\lambda$ $\scriptstyle=(r+\mu)\lambda-\eta$ switching via $c-\lambda$ and $\lambda(t,0)-k$ Formal adjoint for Model E $\scriptstyle\partial_{t}\lambda+\partial_{a}\lambda=$ $\scriptstyle(r+\mu(E,a)+w)\lambda$ $\scriptstyle-cw+\int_{0}^{A}\mu_{E}(E,s)$ $\scriptstyle\times x(t,s)\lambda(t,s)\,ds$ nonlocal coupling through $E(t)$ Main message of the paper The harvesting mechanism is not a cosmetic modeling choice: rate-control $\Rightarrow$ additive / affine / local structure; effort-control $\Rightarrow$ multiplicative / nonlinear / nonlocal structure. This distinction governs persistence, optimality conditions, stationary formulas, and the bioeconomic interpretation.
Figure 1: Mechanism map of the paper. Starting from a unified age-structured transport framework, the analysis splits into rate-control and effort-control harvesting. The former yields a local Pontryagin system; the latter generates a nonlinear nonlocal adjoint through aggregate population dependence.
2 Literature review
The present paper lies at the intersection of five closely related strands of research: age-structured population dynamics, steady-state analysis, optimal control of structured populations, infinite-horizon maximum principles, and bioeconomic harvesting models.
2.1 Age-structured population dynamics
The continuous-time theory of age-structured populations begins with the McKendrick–von Foerster equation, which has become a standard framework for demographic and ecological dynamics [mkendrick_applications_1925, von_foerster_remarks_1959]. A systematic mathematical treatment of nonlinear age-dependent population models is developed by Webb [webb_semigroup_1984, webb_theory_1985], while Iannelli [iannelli_mathematical_1995] and Iannelli and Milner [iannelli_basic_2017] provide rigorous introductions to the analytic and numerical theory, respectively. The broader structured-population perspective, including age and physiological structure, is represented by Metz et al. [metz_dynamics_1986] and Cushing [cushing_introduction_1998]. The semigroup and abstract-evolution approaches may be found in [engel_one-parameter_2000, banasiak_perturbations_2006, magal_theory_2018]. More recent developments on structured models with nonlocal feedback, stability, and lineage-type organization may be found in [wang_damage-structured_2026, liang_global_2025], where feedback-regulated structured dynamics are analyzed in biologically distinct settings.
2.2 Steady states analysis
The analysis of steady states in age-structured models goes back to the nonlinear framework of Gurtin and MacCamy [gurtin_non-linear_1974, gurtin_simple_1979] and the implicit function methods of Cushing [cushing_existence_1984]. Scalar and operator-theoretic reductions of steady-state problems were developed by Prüss [pruss_equilibrium_1981, pruss_stability_1983] and by Busenberg et al. [busenberg_endemic_1988]. A general formulation of nonlinear structured steady states in terms of consistency relations and next-generation operators was given by Diekmann et al. [diekmann_steady-state_2003]. The corresponding threshold phenomena, organized around reproduction numbers, persistence, and extinction, is developed in the work of Thieme [thieme_mathematics_2018] and Inaba [inaba_age-structured_2006]; see also [walker_positive_2009, wang_algebraicspectral_2026, yu_dedifferentiation_2026] for related analyses of threshold conditions and stability.
2.3 Optimal control of age-structured populations
Optimal control for age-structured systems has been studied extensively in both finite- and infinite-horizon settings. Early Pontryagin-type formulations for age-dependent controls were developed by Brokate [brokate_pontryagins_1985], while Barbu and Iannelli [barbu_optimal_1999] established variational formulations and existence results. A comprehensive treatment is given in the monograph of Aniţa [anita_analysis_2000]; see also [anita_optimal_1998]. On the harvesting side, structural properties of optimal policies have been investigated in [hritonenko_structure_2007], while more general control-theoretic expositions may be found in [veliov_optimal_2008]. Furthermore, for stochastic ecological systems, recent analytical frameworks for well-posedness and asymptotic bounds have been developed in [wang_analysis_2025], suggesting a possible route toward stochastic extensions of structured harvesting models.
2.4 Infinite-horizon optimal control and transversality
Infinite-horizon optimal control problems require additional care because the adjoint system must be complemented by a transversality condition at infinity. Classical references include Arrow and Kurz [arrow_optimal_1969] and Carlson et al. [carlson_infinite_1991]. For general infinite-horizon maximum principles, a central role is played by the work of Aseev and Kryazhimskii [aseev_pontryagin_2004], the survey of Aseev et al. [aseev_infinite-horizon_2012], and the refinements of Aseev and Veliov [aseev_maximum_2015]. In the age-structured setting, the interaction between transport dynamics, discounting, and the adjoint system creates additional analytical difficulties; see in particular the discussion of infinite-horizon age-structured optimal control in [aseev_infinite-horizon_2012].
2.5 Rate-control, effort-control, and bioeconomic structure
From the bioeconomic point of view, the distinction between direct harvesting rates and effort-based mortality adjustment is classical. Clark [clark_economics_1973] established the basic economic framework for renewable-resource management, while Getz and Haight [getz_population_1989] and Reed [reed_optimal_1979] analyzed related harvesting mechanisms in demographic and stock-recruitment models. In age-structured settings, the difference is mathematically significant: rate-control enters additively and preserves an affine transport structure, whereas effort-control enters multiplicatively and leads to a bilinear coupling between control and state. When mortality also depends on the aggregate population size, the adjoint equation acquires a nonlocal term coupling all ages through the total stock. Related density-dependent and structured formulations appear in [boucekkine_optimal_2013].
2.6 Position of the present paper
The present paper uses the McKendrick–von Foerster framework as its basic setting, adopts an infinite-horizon variational argument, and places rate-control and effort-control formulations in direct comparison. Its main purpose is not to survey these literature exhaustively, but rather to connect specific issues within a single age-structured framework: the necessary optimality system for the infinite-horizon harvesting problem and the formal emergence of a nonlocal adjoint term in the effort-control case.
3 Model and assumptions
We consider age-structured population models on the age interval $[0,A]$, where $A\in(0,\infty)$ denotes the maximal age and $t\geq 0$ is time. The state variable $x(t,a)\geq 0$ represents the population density of individuals of age $a$ at time $t$. Throughout the paper, $a$ denotes age and $t$ denotes time. The boundary input $p(t)\geq 0$ describes the inflow of newborn or newly introduced individuals at age $a=0$.
Our aim is to compare two harvesting mechanisms within a common age-structured framework. The first one is a rate-control model, in which harvesting acts as a direct removal term. The second one is an effort-control model, in which harvesting enters as an additional mortality intensity and may depend on the aggregate population size. The rate-control model will serve as the main setting for the infinite-horizon maximum principle, while the effort-control model will be used to identify the structural differences induced by multiplicative control and nonlocal population dependence.
3.1 The two model classes
Model R (rate-control).
The age-structured population dynamics with direct harvesting rate $u$ is given by
$\partial_{t}x(t,a)+\partial_{a}x(t,a)$ $=-\mu(t,a)\,x(t,a)-u(t,a),$ $t>0,\;a\in(0,A),$ (1)
$x(t,0)$ $=p(t),$ $t\geq 0,$ (2)
$x(0,a)$ $=x_{0}(a),$ $a\in[0,A].$ (3)
Here $\mu(t,a)\geq 0$ is the natural mortality rate, $u(t,a)\geq 0$ is the harvesting rate applied directly to individuals of age $a$, and $p(t)$ is the boundary recruitment or stocking control.
Model E (effort-control).
The age-structured population dynamics with multiplicative harvesting effort $w$ is
$\partial_{t}x(t,a)+\partial_{a}x(t,a)$ $=-\bigl(\mu(E(t),a)+w(t,a)\bigr)x(t,a),$ $t>0,\;a\in(0,A),$ (4)
$E(t)$ $=\int_{0}^{A}x(t,a)\,da,$ $t\geq 0,$ (5)
$x(t,0)$ $=p(t),$ $t\geq 0,$ (6)
$x(0,a)$ $=x_{0}(a),$ $a\in[0,A].$ (7)
In this formulation, $w(t,a)\geq 0$ is the harvesting effort, while the mortality coefficient $\mu(E(t),a)$ depends on the aggregate population size $E(t)$ and hence introduces a nonlocal dependence into the state equation. Figure 2 illustrates the evolution of the age-structured population along characteristics in both models.
Figure 2: Evolution of the age-structured population along characteristic curves in the $(t,a)$ plane. The state variable $x(t,a)$ propagates along the characteristics $da/dt=1$. The boundary inflow control $p(t)$ is applied at $a=0$, and the initial profile $x_{0}(a)$ is given at $t=0$. The vertical shaded region illustrates the nonlocal aggregate population size $E(t^{*})$, obtained by integrating the population density over all ages at a fixed time $t^{*}$.
The control variables are the distributed harvesting control and the boundary inflow. For Model R, the admissible controls are pairs $(u,p)$ satisfying
$0\leq u(t,a)\leq u_{\max},\qquad 0\leq p(t)\leq p_{\max},\qquad\text{for a.e. }(t,a)\in[0,\infty)\times[0,A],$ (8)
where $u_{\max},p_{\max}\in(0,\infty)$ are fixed constants. For Model E, the admissible controls are pairs $(w,p)$ satisfying
$0\leq w(t,a)\leq w_{\max},\qquad 0\leq p(t)\leq p_{\max},\qquad\text{for a.e. }(t,a)\in[0,\infty)\times[0,A],$ (9)
with $w_{\max},p_{\max}\in(0,\infty)$ fixed. In both model classes, the state is required to remain nonnegative:
$x(t,a)\geq 0\qquad\text{for a.e. }(t,a)\in[0,\infty)\times[0,A].$ (10)
Accordingly, we define the admissible sets
[…]
Throughout the paper we impose the following standing assumptions.
  1. (A1)
    The maximal age $A$ is finite, with $A\in(0,\infty)$.
  2. (A2)
    The initial datum satisfies
    $x_{0}\in L^{1}(0,A),\qquad x_{0}(a)\geq 0\ \text{for a.e. }a\in(0,A).$
  3. (A3)
    In Model R, the natural mortality coefficient
    $\mu:[0,\infty)\times[0,A]\to[0,\infty)$
    is bounded and measurable.
  4. (A4)
    In Model E, the density-dependent mortality coefficient
    $\mu:[0,\infty)\times[0,A]\to[0,\infty),\qquad(E,a)\mapsto\mu(E,a),$
    is bounded on bounded subsets of $0,\infty)\times[0,A]$ and measurable in $a$. When variational arguments are needed for the formal adjoint calculation, we additionally assume that $\mu(\cdot,a)$ is continuously differentiable in $E$ for a.e. $a\in[0,A]$.
  5. (A5)
    The controls belong to the admissible classes $\mathcal{U}_{R}$ or $\mathcal{U}_{E}$, respectively, and the boundary inflow $p$ is nonnegative.
  6. (A6)
    All state trajectories considered in the paper satisfy the nonnegativity constraint (10).
Under these assumptions, each model is understood in the usual mild sense along characteristics. In particular, for every admissible control pair, the corresponding state equation is a first-order transport equation with inflow boundary condition at $a=0$. For the rate-control model this leads to a linear transport problem, while the effort-control model is nonlinear because of the aggregate term $E(t)$ and the multiplicative control structure. These two features are precisely what distinguish the mathematical structure of the two harvesting mechanisms.
The two model classes play different roles in the sequel. For Model R, we study both steady states and an infinite-horizon discounted harvesting problem. In particular, the rate-control framework is used to derive necessary optimality conditions of Pontryagin type. For Model E, our purpose is more limited. We use the effort-control formulation to exhibit how multiplicative harvesting and aggregate dependence modify the structure of the state equation, the objective functional, and the corresponding adjoint system. This will make clear why the effort-control case naturally generates a nonlocal term in the formal dual equation.
To avoid any ambiguity in notation, we summarize the symbols used throughout the paper:
$t\;\text{: time},\qquad a\;\text{: age},\qquad x(t,a)\;\text{: age density},\qquad p(t)\;\text{: boundary inflow},$
[…]
This set of unified notation will be used consistently in all subsequent sections.
4 Infinite-horizon optimal control for the rate-control model
In this section we consider the infinite-horizon harvesting problem associated with the rate-control model. Our goal is to derive a set of first-order necessary conditions in the spirit of Pontryagin’s maximum principle. Since the present paper focuses on the structural form of the optimality system, we formulate the result under the standing assumptions of Section 3 and, in addition, under a regularity framework ensuring that the variational calculations below are justified.
Now we have the optimal control problem. We consider the following infinite-horizon discounted optimization problem.
Problem 4.1.
Maximize
$J(u,p,x)=\int_{0}^{\infty}e^{-rt}\left(\int_{0}^{A}c(t,a)\,u(t,a)\,da-k(t)\,p(t)\right)\,dt$ (13)
over admissible controls $(u,p)\in\mathcal{U}_{R}$ and corresponding states $x$, subject to
$\partial_{t}x(t,a)+\partial_{a}x(t,a)$ $=-\mu(t,a)x(t,a)-u(t,a),$ $t>0,\ a\in(0,A),$ (14)
$x(t,0)$ $=p(t),$ $t\geq 0,$ (15)
$x(0,a)$ $=x_{0}(a),$ $a\in[0,A],$ (16)
$x(t,a)$ $\geq 0$ $\text{for a.e. }(t,a)\in[0,\infty)\times[0,A].$ (17)
Here $r>0$ is the discount rate, $c(t,a)$ is the unit harvesting value, and $k(t)$ is the boundary inflow cost.
In order to interpret the necessary conditions below as a justified first-order optimality system rather than as a purely heuristic calculation, we impose the following assumptions.
  1. (R1)
    Data and admissible controls. The coefficients satisfy
    $c,\mu\in L^{\infty}\bigl((0,\infty)\times(0,A)\bigr),\qquad k\in L^{\infty}(0,\infty),\qquad r>0,$
    and the initial datum satisfies
    $x_{0}\in L^{1}(0,A),\qquad x_{0}(a)\geq 0\ \text{ a.e. }a\in(0,A).$
    The admissible controls satisfy
    $u\in L^{\infty}\bigl((0,\infty)\times(0,A)\bigr),\qquad p\in L^{\infty}(0,\infty),$
    together with the box constraints
    $0\leq u(t,a)\leq u_{\max},\qquad 0\leq p(t)\leq p_{\max}$
    for a.e. $(t,a)$ and a.e. $t$, respectively.
  2. (R2)
    Well-posedness of the state equation. For every admissible pair $(u,p)$, the state equation (14)–(16) admits a unique mild solution
    $x=S(u,p)\in L^{1}_{\mathrm{loc}}\bigl([0,\infty);L^{1}(0,A)\bigr),$
    and this solution satisfies $x(t,a)\geq 0$ for a.e. $(t,a)$.
  3. (R3)
    Differentiability of the control-to-state map. The control-to-state map
    $S:\mathcal{U}_{R}\to L^{1}_{\mathrm{loc}}\bigl([0,\infty);L^{1}(0,A)\bigr),\qquad(u,p)\mapsto x,$
    is Gâteaux differentiable along admissible directions, and for every admissible variation $(\delta u,\delta p)$ the derivative
    $\delta x=DS(u,p)[\delta u,\delta p]$
    is the unique mild solution of the linearized system associated with (14)–(16).
  4. (R4)
    Integrability of the objective functional. For every admissible pair $(u,p)$, the discounted objective is well defined:
    $\int_{0}^{\infty}e^{-rt}\left(\int_{0}^{A}|c(t,a)u(t,a)|\,da+|k(t)p(t)|\right)\,dt<\infty.$
  5. (R5)
    Regularity of adjoint and multiplier. The adjoint variable $\lambda$ and the state-constraint multiplier $\eta$ belong to classes for which all terms in the Lagrangian and in the integration-by-parts formulas are meaningful, and such that
    $e^{-rt}\lambda\in W^{1,1}_{\mathrm{loc}}\bigl((0,\infty)\times(0,A)\bigr),\qquad\eta\in L^{1}_{\mathrm{loc}}\bigl((0,\infty)\times(0,A)\bigr),\qquad\eta\geq 0.$
  6. (R6)
    Justification of integration by parts and transversality. The products
    $e^{-rt}\lambda\,\delta x,\qquad e^{-rt}\lambda\,\partial_{t}\delta x,\qquad e^{-rt}\lambda\,\partial_{a}\delta x$
    are integrable on finite cylinders, and the integration-by-parts identities used are valid. Moreover, for every admissible first variation $\delta x$, the boundary pairing at infinity vanishes:
    $\lim_{T\to\infty}\int_{0}^{A}e^{-rT}\lambda(T,a)\delta x(T,a)\,da=0.$
  7. (R7)
    Existence of a KKT multiplier for the state constraint. The nonnegativity constraint on the state is qualified in a form that guarantees the existence of a multiplier $\eta\geq 0$ satisfying the complementary slackness relation
    $\eta(t,a)x^{*}(t,a)=0\qquad\text{for a.e. }(t,a)\in[0,\infty)\times[0,A].$
Under assumptions (R1)–(R7), the variational calculation leading to the necessary conditions below is understood as a justified first-order derivation of the optimality system for Problem 4.1.
To derive the optimality system, we introduce a current-value adjoint variable $\lambda(t,a)$ for the state equation and a current-value multiplier $\eta(t,a)\geq 0$ for the state constraint $x(t,a)\geq 0$. Formally, consider the Lagrangian
$\mathcal{L}(u,p,x;\lambda,\eta)$ $=\int_{0}^{\infty}e^{-rt}\left(\int_{0}^{A}c(t,a)\,u(t,a)\,da-k(t)\,p(t)\right)\,dt$
$\quad-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\Bigl(\partial_{t}x(t,a)+\partial_{a}x(t,a)+\mu(t,a)x(t,a)+u(t,a)\Bigr)\,da\,dt$
$\quad+\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\eta(t,a)\,x(t,a)\,da\,dt.$ (18)
Let $(u^{*},p^{*},x^{*})$ be an optimal triple. For admissible perturbations $(\delta u,\delta p)$, define
$u^{\varepsilon}:=u^{*}+\varepsilon\,\delta u,\qquad p^{\varepsilon}:=p^{*}+\varepsilon\,\delta p,$
for $|\varepsilon|$ sufficiently small, chosen so that feasibility is preserved. Let $x^{\varepsilon}$ denote the corresponding perturbed state, and let
$\delta x:=\left.\frac{d}{d\varepsilon}x^{\varepsilon}\right|_{\varepsilon=0}$
be the induced first variation of the state. By differentiating the state equation with respect to $\varepsilon$, one obtains the linearized system
$\partial_{t}\delta x(t,a)+\partial_{a}\delta x(t,a)$ $=-\mu(t,a)\delta x(t,a)-\delta u(t,a),$ (19)
$\delta x(t,0)$ $=\delta p(t),\qquad\delta x(0,a)=0.$ (20)
We now evaluate the first variation of the current-value Lagrangian $\mathcal{L}$ in (18). Differentiation at $\varepsilon=0$ yields
$\delta\mathcal{L}$ $=\int_{0}^{\infty}e^{-rt}\left(\int_{0}^{A}c(t,a)\,\delta u(t,a)\,da-k(t)\,\delta p(t)\right)\,dt$
$\quad-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\Bigl(\partial_{t}\delta x(t,a)+\partial_{a}\delta x(t,a)+\mu(t,a)\delta x(t,a)+\delta u(t,a)\Bigr)\,da\,dt$
$\quad+\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\eta(t,a)\,\delta x(t,a)\,da\,dt.$ (21)
After collecting the $\delta u$-terms, this becomes
$\delta\mathcal{L}$ $=\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\bigl(c(t,a)-\lambda(t,a)\bigr)\delta u(t,a)\,da\,dt-\int_{0}^{\infty}e^{-rt}k(t)\,\delta p(t)\,dt$
$\quad-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\bigl(\partial_{t}\delta x(t,a)+\partial_{a}\delta x(t,a)\bigr)\,da\,dt$
$\quad+\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\bigl(-\mu(t,a)\lambda(t,a)+\eta(t,a)\bigr)\delta x(t,a)\,da\,dt.$ (22)
We next transfer derivatives from $\delta x$ to the adjoint variable by integration by parts. For the time derivative term, one has
$-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\,\partial_{t}\delta x(t,a)\,da\,dt$ $=-\lim_{T\to\infty}\int_{0}^{A}e^{-rT}\lambda(T,a)\delta x(T,a)\,da$
$\quad+\int_{0}^{A}\lambda(0,a)\delta x(0,a)\,da$
$\quad+\int_{0}^{\infty}\int_{0}^{A}\partial_{t}\bigl(e^{-rt}\lambda(t,a)\bigr)\delta x(t,a)\,da\,dt.$ (23)
Since $\delta x(0,a)=0$, the term at $t=0$ vanishes. Moreover,
$\partial_{t}\bigl(e^{-rt}\lambda(t,a)\bigr)=e^{-rt}\bigl(\partial_{t}\lambda(t,a)-r\lambda(t,a)\bigr).$
Similarly, for the age derivative term,
$-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\,\partial_{a}\delta x(t,a)\,da\,dt$ $=-\int_{0}^{\infty}e^{-rt}\lambda(t,A)\delta x(t,A)\,dt$
$\quad+\int_{0}^{\infty}e^{-rt}\lambda(t,0)\delta x(t,0)\,dt$
$\quad+\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\partial_{a}\lambda(t,a)\delta x(t,a)\,da\,dt.$ (24)
Using $\delta x(t,0)=\delta p(t)$, substitution of (23) and (24) into (22) yields
$\delta\mathcal{L}$ $=\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\bigl(c(t,a)-\lambda(t,a)\bigr)\delta u(t,a)\,da\,dt$
$\quad+\int_{0}^{\infty}e^{-rt}\bigl(\lambda(t,0)-k(t)\bigr)\delta p(t)\,dt$
$\quad+\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\Bigl(\partial_{t}\lambda(t,a)+\partial_{a}\lambda(t,a)-\bigl(r+\mu(t,a)\bigr)\lambda(t,a)+\eta(t,a)\Bigr)\delta x(t,a)\,da\,dt$
$\quad-\int_{0}^{\infty}e^{-rt}\lambda(t,A)\delta x(t,A)\,dt-\lim_{T\to\infty}\int_{0}^{A}e^{-rT}\lambda(T,a)\delta x(T,a)\,da.$ (25)
The coefficient of the interior state variation vanishes provided the adjoint satisfies
$\partial_{t}\lambda(t,a)+\partial_{a}\lambda(t,a)=\bigl(r+\mu(t,a)\bigr)\lambda(t,a)-\eta(t,a),\qquad t>0,\ a\in(0,A).$
The boundary contribution at $a=A$ is removed by imposing
$\lambda(t,A)=0,\qquad t\geq 0.$
Finally, the boundary term at infinity disappears under the transversality condition:
$\lim_{T\to\infty}\int_{0}^{A}e^{-rT}\lambda(T,a)\delta x(T,a)\,da=0$
for every admissible first variation $\delta x$. Under these choices, the first variation reduces to
$\delta\mathcal{L}=\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\bigl(c(t,a)-\lambda(t,a)\bigr)\delta u(t,a)\,da\,dt+\int_{0}^{\infty}e^{-rt}\bigl(\lambda(t,0)-k(t)\bigr)\delta p(t)\,dt.$ (26)
This identifies the switching gradients
$\frac{\partial I}{\partial u}(t,a)=e^{-rt}\bigl(c(t,a)-\lambda(t,a)\bigr),\qquad\frac{\partial I}{\partial p}(t)=e^{-rt}\bigl(\lambda(t,0)-k(t)\bigr).$
We now collect the preceding calculations into the main result of this section.
Theorem 4.2 (Necessary conditions).
Let $(u^{*},p^{*},x^{*})$ be an optimal solution of Problem 4.1. Assume that (R1)(R7) hold. Then there exist a current-value adjoint variable
$\lambda:[0,\infty)\times[0,A]\to\mathbb{R}$
and a current-value state-constraint multiplier
$\eta:[0,\infty)\times[0,A]\to[0,\infty)$
such that the following conditions hold.
  1. (i)
    Adjoint equation. The adjoint variable satisfies
    $\partial_{t}\lambda(t,a)+\partial_{a}\lambda(t,a)=\bigl(r+\mu(t,a)\bigr)\lambda(t,a)-\eta(t,a),\qquad t>0,\ a\in(0,A).$ (27)
  2. (ii)
    Terminal-age condition and transversality. One has
    $\lambda(t,A)=0\qquad\text{for }t\geq 0,$ (28)
    and
    $\lim_{T\to\infty}\int_{0}^{A}e^{-rT}\lambda(T,a)\delta x(T,a)\,da=0$ (29)
    for every admissible first variation $\delta x$.
  3. (iii)
    Switching conditions for the distributed control $u$. Define
    $\frac{\partial I}{\partial u}(t,a):=e^{-rt}\bigl(c(t,a)-\lambda(t,a)\bigr).$ (30)
    Then
    $\frac{\partial I}{\partial u}(t,a)\leq 0$ $\qquad\text{at points where }u^{*}(t,a)=0,$ (31)
    $\frac{\partial I}{\partial u}(t,a)=0$ $\qquad\text{at points where }0<u^{*}(t,a)<u_{\max},$ (32)
    $\frac{\partial I}{\partial u}(t,a)\geq 0$ $\qquad\text{at points where }u^{*}(t,a)=u_{\max}.$ (33)
  4. (iv)
    Switching conditions for the boundary control $p$. Define
    $\frac{\partial I}{\partial p}(t):=e^{-rt}\bigl(\lambda(t,0)-k(t)\bigr).$ (34)
    Then
    $\frac{\partial I}{\partial p}(t)\leq 0$ $\qquad\text{at points where }p^{*}(t)=0,$ (35)
    $\frac{\partial I}{\partial p}(t)=0$ $\qquad\text{at points where }0<p^{*}(t)<p_{\max},$ (36)
    $\frac{\partial I}{\partial p}(t)\geq 0$ $\qquad\text{at points where }p^{*}(t)=p_{\max}.$ (37)
  5. (v)
    Complementary slackness for the state constraint. The multiplier $\eta$ satisfies
    $\eta(t,a)\geq 0,\qquad x^{*}(t,a)\geq 0,\qquad\eta(t,a)x^{*}(t,a)=0\qquad\text{for a.e. }(t,a)\in[0,\infty)\times[0,A].$ (38)
Proof.
The necessary conditions are obtained from the first variation of the current-value Lagrangian together with the normal-cone conditions associated with the box constraints and the complementary slackness relation for the state inequality constraint.
We now exploit the first-order optimality condition $\delta\mathcal{L}\leq 0$ for every feasible direction at the optimal pair. Consider first the distributed control $u$.
If $0<u^{*}(t,a)<u_{\max}$, then both positive and negative local perturbations of $u$ are admissible. Hence the coefficient of $\delta u$ in (26) must vanish, which gives
$\frac{\partial I}{\partial u}(t,a)=0.$
If $u^{*}(t,a)=0$, then only nonnegative local perturbations are admissible, and therefore
$\frac{\partial I}{\partial u}(t,a)\leq 0.$
If $u^{*}(t,a)=u_{\max}$, then only nonpositive local perturbations are admissible, and thus
$\frac{\partial I}{\partial u}(t,a)\geq 0.$
This proves (31)–(33).
The argument for the boundary control $p$ is identical. If $0<p^{*}(t)<p_{\max}$, both signs of $\delta p(t)$ are admissible and therefore
$\frac{\partial I}{\partial p}(t)=0.$
If $p^{*}(t)=0$, only nonnegative perturbations are admissible, which yields
$\frac{\partial I}{\partial p}(t)\leq 0.$
If $p^{*}(t)=p_{\max}$, only nonpositive perturbations are admissible, and hence
$\frac{\partial I}{\partial p}(t)\geq 0.$
This proves (35)–(37).
It remains to treat the state inequality constraint. Since the condition $x^{*}(t,a)\geq 0$ is enforced through the multiplier $\eta(t,a)\geq 0$, the Karush–Kuhn–Tucker conditions yield
$\eta(t,a)\geq 0,\qquad x^{*}(t,a)\geq 0,\qquad\eta(t,a)x^{*}(t,a)=0$
for a.e. $(t,a)\in[0,\infty)\times[0,A]$, which is precisely (38). This completes the proof. ∎
Remark 4.3 (Scope of the theorem).
Theorem 4.2 is stated as a necessary optimality system for the rate-control problem under the regularity and well-posedness assumptions needed to justify the variational calculation. On an unbounded time horizon, the role of the adjoint variable and of the transversality condition at infinity requires particular care in a fully rigorous functional-analytic treatment. In the present paper, the purpose of this section is to identify the structural form of the adjoint system, the switching laws, and the complementary slackness relation that will later be compared with the effort-control formulation.
5 Effort-control model and the nonlocal adjoint term
We now turn to the effort-control formulation introduced in Section 3. In contrast to the rate-control model, the state equation is nonlinear because the mortality coefficient depends on the aggregate population size
$E(t)=\int_{0}^{A}x(t,a)\,da,$
and the control enters multiplicatively through the term $w(t,a)x(t,a)$. For this reason, and in order not to overstate the level of rigor at this stage, we restrict ourselves here to a formal derivation of the associated adjoint system under sufficient smoothness assumptions.
For the effort-control problem, consider the infinite-horizon discounted functional
$J_{E}(w,p,x)=\int_{0}^{\infty}e^{-rt}\left(\int_{0}^{A}c(t,a)\,w(t,a)x(t,a)\,da-k(t)\,p(t)\right)\,dt,$ (39)
subject to the nonlinear age-structured dynamics
$\partial_{t}x(t,a)+\partial_{a}x(t,a)$ $=-\bigl(\mu(E(t),a)+w(t,a)\bigr)x(t,a),$ $t>0,\ a\in(0,A),$ (40)
$E(t)$ $=\int_{0}^{A}x(t,a)\,da,$ $t\geq 0,$ (41)
$x(t,0)$ $=p(t),$ $t\geq 0,$ (42)
$x(0,a)$ $=x_{0}(a),$ $a\in[0,A].$ (43)
The purpose of this section is not to establish a complete Pontryagin principle for (39)–(43), but rather to identify the form of the nonlocal term that appears in the dual equation when the mortality rate depends on the aggregate state variable $E(t)$.
Assume for the moment that all data are sufficiently smooth and that the state, control, and adjoint variables are regular enough for differentiation under the integral sign and integration by parts to be justified. Introduce a current-value adjoint variable $\lambda(t,a)$ for the state equation and write the current-value Lagrangian in the form
$\mathcal{L}_{E}(w,p,x;\lambda)$ $=\int_{0}^{\infty}e^{-rt}\left(\int_{0}^{A}c(t,a)\,w(t,a)x(t,a)\,da-k(t)\,p(t)\right)\,dt$
$\quad-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\Bigl(\partial_{t}x(t,a)+\partial_{a}x(t,a)+\bigl(\mu(E(t),a)+w(t,a)\bigr)x(t,a)\Bigr)\,da\,dt.$ (44)
Let $(w^{*},p^{*},x^{*})$ be an admissible triple and let $\delta x$ be the first variation of the state induced by admissible perturbations. The terms involving $\delta x$ arise from three sources:
  1. (i)
    the payoff density $c(t,a)w(t,a)x(t,a)$;
  2. (ii)
    the local term $\bigl(\mu(E(t),a)+w(t,a)\bigr)x(t,a)$ in the state equation;
  3. (iii)
    the dependence of $\mu(E(t),a)$ on the aggregate state
    $E(t)=\int_{0}^{A}x(t,s)\,ds.$
The first two sources give the expected local contributions. The third one generates an additional nonlocal term. Indeed, differentiating $\mu(E(t),a)x(t,a)$ with respect to the state yields
$\delta\!\bigl(\mu(E(t),a)x(t,a)\bigr)=\mu(E(t),a)\,\delta x(t,a)+\mu_{E}(E(t),a)\,x(t,a)\,\delta E(t),$
where
$\delta E(t)=\int_{0}^{A}\delta x(t,s)\,ds.$
Substituting this into the first variation of (44), the nonlocal contribution becomes
$-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\mu_{E}(E(t),a)x(t,a)\,\delta E(t)\,da\,dt$
$\qquad=-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\lambda(t,a)\mu_{E}(E(t),a)x(t,a)\left(\int_{0}^{A}\delta x(t,s)\,ds\right)\,da\,dt$
$\qquad=-\int_{0}^{\infty}\int_{0}^{A}e^{-rt}\left(\int_{0}^{A}\mu_{E}(E(t),\sigma)x(t,\sigma)\lambda(t,\sigma)\,d\sigma\right)\delta x(t,s)\,ds\,dt.$ (45)
Therefore, this contributes the coefficient
$\int_{0}^{A}\mu_{E}(E(t),s)x(t,s)\lambda(t,s)\,ds$
to the interior $\delta x$-variation.
The local terms are handled exactly as in the rate-control case: integrating by parts in $t$ and $a$ transfers the derivatives from $\delta x$ onto the adjoint, while the discount factor produces the term $r\lambda$. The contribution of the running payoff $c(t,a)w(t,a)x(t,a)$ yields the local source term c(t,a)w(t,a)-c(t,a)w(t,a) in the adjoint equation, and the multiplicative mortality term $w(t,a)x(t,a)$ contributes $w(t,a)\lambda(t,a)$.
These formal calculations lead to the following statement.
Proposition 5.1 (Formal adjoint system for the effort-control problem).
Assume that the coefficients, controls, and state are sufficiently smooth, that $\mu(E,a)$ is continuously differentiable with respect to $E$, and that all integration-by-parts steps are justified. Then the current-value adjoint variable associated with (39)–(43) formally satisfies
$\partial_{t}\lambda(t,a)+\partial_{a}\lambda(t,a)$ $=\bigl(r+\mu(E(t),a)+w(t,a)\bigr)\lambda(t,a)-c(t,a)w(t,a)$
$\quad+\int_{0}^{A}\mu_{E}(E(t),s)x(t,s)\lambda(t,s)\,ds,\qquad t>0,\ a\in(0,A).$ (46)
The corresponding terminal-age and transversality conditions are formally given by
$\begin{aligned} &\qquad\qquad\lambda(t,A)=0\qquad\text{for }t\geq 0,\\[4.0pt] &\lim_{T\to\infty}\int_{0}^{A}e^{-rT}\lambda(T,a)\delta x(T,a)\,da=0\quad\text{for every admissible first variation }\delta x.\end{aligned}$ (47)
Remark 5.2.
The structural contrast with the rate-control model is now transparent. In the rate-control case, the adjoint equation is local in the age variable and involves only the state-constraint multiplier as a source term. In the effort-control case, the dependence of $\mu$ on the aggregate population size produces the additional nonlocal term
$\int_{0}^{A}\mu_{E}(E(t),s)x(t,s)\lambda(t,s)\,ds,$
which couples all ages through the total population level $E(t)$. This is the main analytical feature distinguishing the effort-control formulation from the rate-control formulation.
6 Autonomous reductions and explicit stationary formulas
This section collects several autonomous reductions and explicit formulas that complement the general theory developed above. Their role is not to introduce new existence theory, but rather to provide computable representations of the state and adjoint variables in the stationary setting.
5Throughout this section we assume that the coefficients are time-independent. For the rate-control model we therefore write
$\mu(t,a)=\mu(a),\qquad c(t,a)=c(a),\qquad k(t)=k,$ (48)
and we seek steady states of the form
$x(t,a)=x(a),\qquad u(t,a)=u(a),\qquad p(t)=p.$ (49)
Similarly, for the effort-control model we assume
$\mu(E(t),a)=\mu(E,a),\qquad c(t,a)=c(a),\qquad k(t)=k,$ (50)
and seek steady states of the form
$x(t,a)=x(a),\qquad w(t,a)=w(a),\qquad p(t)=p,\qquad E(t)=E.$ (51)
Under these autonomous assumptions, the transport equations reduce to first-order ordinary differential equations in the age variable.
Proposition 6.1 (Steady-state ODE for the rate-control model).
Under (48)–(49), the rate-control state equation
$\partial_{t}x(t,a)+\partial_{a}x(t,a)=-\mu(t,a)x(t,a)-u(t,a),\qquad x(t,0)=p(t),$
reduces to
$x^{\prime}(a)=-\mu(a)x(a)-u(a),\qquad x(0)=p.$ (52)
Proof.
Since $x(t,a)=x(a)$ is independent of $t$, one has $\partial_{t}x(t,a)=0$. Substituting (48)–(49) into the state equation yields (52). The boundary condition $x(t,0)=p(t)$ reduces to $x(0)=p$. ∎
Proposition 6.2 (Steady-state ODE for the effort-control model).
Under (50)–(51), the effort-control state equation
$\partial_{t}x(t,a)+\partial_{a}x(t,a)=-\bigl(\mu(E(t),a)+w(t,a)\bigr)x(t,a),\qquad x(t,0)=p(t),$
reduces to
$x^{\prime}(a)=-\bigl(\mu(E,a)+w(a)\bigr)x(a),\qquad x(0)=p.$ (53)
Proof.
Under the steady-state ansatz, $\partial_{t}x(t,a)=0$, and direct substitution into the state equation gives (53). The boundary condition again reduces to $x(0)=p$. ∎
The two autonomous state equations admit explicit integral representations.
Proposition 6.3 (Explicit stationary profile for the rate-control model).
The unique solution of (52) is
$x(a)=p\exp\!\left(-\int_{0}^{a}\mu(\xi)\,d\xi\right)-\int_{0}^{a}\exp\!\left(-\int_{s}^{a}\mu(\xi)\,d\xi\right)u(s)\,ds,\qquad 0\leq a\leq A.$ (54)
Proof.
Equation (52) may be rewritten as
$x^{\prime}(a)+\mu(a)x(a)=-u(a).$
Introduce the integrating factor
$M(a):=\exp\!\left(\int_{0}^{a}\mu(\xi)\,d\xi\right).$
Then
$\frac{d}{da}\bigl(M(a)x(a)\bigr)=-M(a)u(a).$
Integrating from 0 to $a$ and using the boundary condition $x(0)=p$ yields
$M(a)x(a)=p-\int_{0}^{a}M(s)u(s)\,ds.$
Dividing by $M(a)$ gives
$x(a)=p\exp\!\left(-\int_{0}^{a}\mu(\xi)\,d\xi\right)-\int_{0}^{a}\exp\!\left(-\int_{s}^{a}\mu(\xi)\,d\xi\right)u(s)\,ds,$
which is (54). ∎
Proposition 6.4 (Explicit stationary profile for the effort-control model).
The unique solution of (53) is
$x(a)=p\exp\!\left(-\int_{0}^{a}\bigl(\mu(E,\xi)+w(\xi)\bigr)\,d\xi\right),\qquad 0\leq a\leq A.$ (55)
Proof.
Equation (53) is separable:
$\frac{x^{\prime}(a)}{x(a)}=-\bigl(\mu(E,a)+w(a)\bigr).$
Integrating from 0 to $a$ and using $x(0)=p$ gives
$\log x(a)-\log p=-\int_{0}^{a}\bigl(\mu(E,\xi)+w(\xi)\bigr)\,d\xi,$
and therefore
$x(a)=p\exp\!\left(-\int_{0}^{a}\bigl(\mu(E,\xi)+w(\xi)\bigr)\,d\xi\right),$
which is (55). ∎
We next record the autonomous form of the adjoint equation for the rate-control model.
Proposition 6.5 (Adjoint ODE and explicit representation).
Assume that the data are autonomous and that the current-value adjoint $\lambda(t,a)$ is stationary, i.e.
$\lambda(t,a)=\lambda(a),\qquad\eta(t,a)=\eta(a).$
Then the adjoint PDE
$\partial_{t}\lambda(t,a)+\partial_{a}\lambda(t,a)=\bigl(r+\mu(t,a)\bigr)\lambda(t,a)-\eta(t,a),\qquad\lambda(t,A)=0,$
reduces to
$\lambda^{\prime}(a)=\bigl(r+\mu(a)\bigr)\lambda(a)-\eta(a),\qquad\lambda(A)=0.$ (56)
Moreover, the unique solution is
$\lambda(a)=\int_{a}^{A}\exp\!\left(-\int_{a}^{s}\bigl(r+\mu(\xi)\bigr)\,d\xi\right)\eta(s)\,ds,\qquad 0\leq a\leq A.$ (57)
Proof.
Under stationarity, $\partial_{t}\lambda(t,a)=0$, and the adjoint PDE reduces to (56). Rewrite the equation as
$\lambda^{\prime}(a)-\bigl(r+\mu(a)\bigr)\lambda(a)=-\eta(a).$
Define the integrating factor
$N(a):=\exp\!\left(-\int_{0}^{a}\bigl(r+\mu(\xi)\bigr)\,d\xi\right).$
Then
$\frac{d}{da}\bigl(N(a)\lambda(a)\bigr)=N(a)\lambda^{\prime}(a)+N^{\prime}(a)\lambda(a)=N(a)\lambda^{\prime}(a)-N(a)\bigl(r+\mu(a)\bigr)\lambda(a)=-N(a)\eta(a).$
Integrating from $a$ to AA and using the terminal condition $\lambda(A)=0$ yields
$N(A)\lambda(A)-N(a)\lambda(a)=-\int_{a}^{A}N(s)\eta(s)\,ds,$
hence
$N(a)\lambda(a)=-\int_{a}^{A}N(s)\eta(s)\,ds.$
Dividing by N(a)-N(a) gives
$\lambda(a)=\int_{a}^{A}\frac{N(s)}{N(a)}\eta(s)\,ds.$
Finally,
$\frac{N(s)}{N(a)}=\exp\!\left(-\int_{0}^{s}\bigl(r+\mu(\xi)\bigr)\,d\xi+\int_{0}^{a}\bigl(r+\mu(\xi)\bigr)\,d\xi\right)=\exp\!\left(-\int_{a}^{s}\bigl(r+\mu(\xi)\bigr)\,d\xi\right),$
and (57) follows. ∎
Figure 3 presents the information flow and system architecture of the infinite-horizon necessary optimality conditions.
Figure 3: Information flow and system architecture of the infinite-horizon necessary optimality conditions. The system illustrates the fundamental coupling between the forward-in-time demographic state dynamics and the backward-in-time adjoint dynamics (originating from the transversality conditions at $T\to\infty$ and $a=A$). The shadow prices $\lambda$ generated by the adjoint system subsequently dictate the optimal distributed and boundary controls via the switching conditions.
The propositions of this section provide explicit stationary representations that can be used either for direct computation or for the qualitative interpretation of the optimality results established earlier. In particular, the difference between the rate-control and effort-control formulations becomes transparent at the level of the stationary profiles: the former leads to an affine Volterra representation (54), whereas the latter yields the purely multiplicative exponential profile (55).
7 Discussion
This paper presents two different formulations: rate-control and effort-control harvesting formulations. For the rate-control formulation, the associated infinite-horizon optimization problem admits a Pontryagin-type necessary optimality system. In current-value form, this system consists of the adjoint transport equation, the terminal-age and transversality conditions. Besides, we study the switching relations for the distributed control and the boundary inflow, and the complementary slackness condition associated with the state constraint $x\geq 0$.
A central contribution of the paper is the clarification of the structural divide between rate-control and effort-control mechanisms. In the rate-control model, harvesting enters additively, which leads to an affine Volterra-type stationary profile and a local adjoint equation. By contrast, effort-control acts multiplicatively. When natural mortality depends on the aggregate population size $E(t)$, the effort-control adjoint acquires the nonlocal coupling term
$\int_{0}^{A}\mu_{E}(E(t),s)x(t,s)\lambda(t,s)\,ds,$
which couples all ages through the total population level. Thus, direct removal and mortality-intensity adjustment are not merely alternative parameterizations of the same model, but generate genuinely different analytical structures.
From a methodological point of view, the paper provides a rigorous variational derivation for the rate-control optimality system, while the effort-control problem is treated at the level of a formal adjoint calculation. This distinction is deliberate: the nonlocal and nonlinear structure of the effort-control dynamics requires a separate well-posedness and differentiability theory.
Natural directions for future work include the dynamical stability of the supercritical positive equilibrium, a fully rigorous functional-analytic treatment of the infinite-horizon transversality condition, and the well-posedness of the nonlinear nonlocal effort-control state equation. First, it would be interesting to connect the present structured-population framework with multiscale hybrid models that combine continuum dynamics with agent-based components [liu_bidirectional_2025, wang_analysis_2025-1]. The hybrid framework will be a natural setting to incorporate more complex environmental factors. Second, it would also be interesting to connect the present age-structured framework with hybrid PDE-based models that combine transport, feedback, and mode-selection mechanisms in multiscale environments, as in recent studies of tumor–stroma patterning and hybrid PDE–ODE mass-transport systems [yu_chemotactic_2026, yu_mode-wise_2026]. A further direction is the incorporation of demographic fluctuations into structured harvesting models. Recent stochastic ecological work has emphasized that covariance-consistent noise and absorbing-boundary formulations can be essential for correctly capturing extinction structure [yu_full-covariance_2026, yu_beyond_2026].
Appendix A Numerical simulations
In this section we illustrate the analytical results of the preceding sections by means of numerical computations. Unless otherwise stated, the simulations use the common parameter set
$A=10,\qquad r=0.05,\qquad\mu(a)=0.01+0.005\,a,$
so that the natural mortality increases linearly with age. Unless otherwise stated, the age interval $[0,A]$ is discretized with $500$ grid points. The accompanying Python code reproduces all figures below after specifying a valid output directory.
A.1 Stationary age profiles under harvesting
Figure 4 compares stationary age profiles for the two harvesting mechanisms. In both cases, the boundary inflow is $p=1$ and harvesting is active on the age window $[3,7]$ with uniform intensity $h=0.08$.
Figure 4: Stationary age profiles. (a) Rate-control: the direct removal $u(a)$ produces a visible change in slope at the harvesting window boundaries. (b) Effort-control: the multiplicative effort $w(a)$ preserves the exponential structure, yielding a smaller baseline population density. The shaded region indicates the age window $[3,7]$ where harvesting is active. Dashed curves show the unharvested baseline.
For the rate-control model, the stationary profile is computed numerically from the steady-state ODE and is consistent with the explicit representation of Proposition 6.3,
$x(a)=p\exp\!\Bigl(-\int_{0}^{a}\mu(\xi)\,d\xi\Bigr)-\int_{0}^{a}\exp\!\Bigl(-\int_{s}^{a}\mu(\xi)\,d\xi\Bigr)u(s)\,ds.$
The resulting profile exhibits a visible change in slope at the boundaries of the harvesting window ($a=3$ and $a=7$), reflecting the additive removal term $u(a)$.
For the effort-control model, the stationary profile is given by (Proposition 6.4)
$x(a)=p\exp\!\Bigl(-\int_{0}^{a}\bigl(\mu(E,\xi)+w(\xi)\bigr)\,d\xi\Bigr),$
where
$E=\int_{0}^{A}x(a)\,da$
is determined self-consistently by fixed-point iteration with density-dependent mortality
$\mu(E,a)=\mu(a)+\alpha E,\qquad\alpha=0.002.$
In contrast to the rate-control case, the profile exhibits a lower baseline population density at the same age, while the difference between the control densities is smaller.
A.2 Time-dependent dynamics
To illustrate the transient behavior of the rate-control model, we solve the PDE (1)–(3) numerically by an upwind finite-difference scheme on a grid with $200$ age points and $400$ time steps over the horizon $[0,20]$. The boundary inflow is periodic:
$p(t)=0.5+0.3\sin(2\pi t/8),$
representing seasonal restocking. Harvesting is ramped up linearly over the interval $[0,5]$ to its target value $u=0.06$ on the age window $[3,7]$.
Figure 5 shows the result. Panel (a) displays the population density $x(t,a)$ as a space–time heatmap, where the characteristic diagonal transport structure is clearly visible. Panel (b) shows age profiles at selected times, illustrating how the initial condition evolves under the combined effect of transport, harvesting, and periodic boundary input. Panel (c) tracks the aggregate population
$E(t)=\int_{0}^{A}x(t,a)\,da,$
which exhibits an initial transient followed by oscillations around a long-run mean, consistent with the periodic forcing at the boundary.
Figure 5: Time-dependent dynamics of the rate-control model with periodic boundary inflow and ramped harvesting. (a) Space–time heatmap of the population density. (b) Age profiles at selected times. (c) Aggregate population $E(t)$ showing an initial transient followed by quasi-periodic oscillations.
A.3 Adjoint variable and switching structure
Figure 6 provides an illustrative stationary example of the adjoint variable and the associated switching law from Theorem 4.2. We prescribe a nontrivial multiplier profile $\eta(a)$ near the maximal age in order to visualize how the shadow value $\lambda(a)$ affects the harvesting decision; specifically, we take $\eta(a)>0$ for $a\geq 8.5$. The stationary adjoint ODE (Proposition 6.5)
$\lambda^{\prime}(a)=\bigl(r+\mu(a)\bigr)\lambda(a)-\eta(a),\qquad\lambda(A)=0,$
is solved backward from $a=A$. The unit harvesting value is chosen as
$c(a)=1+0.5\sin\!\bigl(\pi(a-2)/6\bigr)\quad\text{for }a\in[2,8],$
and
$c(a)=0.2\quad\text{otherwise}.$
Figure 6: Illustrative stationary adjoint and switching structure for the rate-control model. (a) Adjoint variable $\lambda(a)$, unit value $c(a)$, cost $k$, and prescribed multiplier $\eta(a)$. (b) Switching function $\sigma_{u}(a)=c(a)-\lambda(a)$. (c) Resulting bang-bang proxy control $u^{*}(a)$.
Panel (a) displays the adjoint $\lambda(a)$ together with the value profile $c(a)$, the boundary inflow cost $k=0.6$, and the prescribed multiplier $\eta(a)$. Panel (b) plots the switching function
$\sigma_{u}(a)=c(a)-\lambda(a).$
Positive values indicate ages at which harvesting is favored, while negative values indicate ages at which conservation is favored. Panel (c) shows the resulting bang-bang proxy control $u^{*}(a)$ obtained by activating harvesting at full intensity ($u_{\max}=0.15$) when the switching function is sufficiently positive and setting it to zero otherwise.
A.4 Rate-control versus effort-control: yield and depletion
Finally, Figure 7 compares the two harvesting mechanisms by sweeping a common harvesting intensity $h$ over the interval $[0,0.5]$.
Figure 7: Comparison of rate-control and effort-control harvesting. (a) Stationary yield as a function of harvesting intensity: the rate-control yield grows rapidly and then levels off under the numerical truncation induced by the state constraint, whereas the effort-control yield is concave throughout. (b) Aggregate population depletion: rate-control depletes the stock more aggressively at equal intensity.
For the rate-control model, we set
$u(a)=h\,\mathbf{1}_{[2,8]}(a),$
and define the stationary yield by
$Y_{R}(h)=\int_{0}^{A}u_{\mathrm{actual}}(a)\,da,$
where $u_{\mathrm{actual}}$ denotes the numerically truncated extraction rate used to enforce the nonnegativity constraint $x(a)\geq 0$.
For the effort-control model, we set
$w(a)=h\,\mathbf{1}_{[2,8]}(a),$
and define
$Y_{E}(h)=\int_{0}^{A}w(a)x(a)\,da,$
where $x(a)$ depends on $h$ through the steady-state equation with density-dependent mortality
$\mu(E,a)=\mu(a)+\alpha E,\qquad\alpha=0.002.$
Panel (a) shows that the rate-control yield initially increases steeply and then levels off once the nonnegativity constraint becomes active through the numerical truncation of the admissible extraction. By contrast, the effort-control yield is concave from the outset, since higher effort simultaneously depletes the stock on which the yield depends. Panel (b) shows the corresponding aggregate population
$E(h)=\int_{0}^{A}x(a)\,da$
under each mechanism. At a given intensity, the rate-control model depletes the stock more aggressively than the effort-control model, which is partially self-regulated by the multiplicative structure $w(a)x(a)$.